OpenStudy (anonymous):

Given that the rate constant for the decomposition of hypothetical compound X from part A is 1.60 M−1⋅min−1, calculate the concentration of X after 13.0min . (min) [X](M) 0 0.467 1 0.267 2 0.187 3 0.144 4 0.117 5 0.099 6 0.085 7 0.075

OpenStudy (aaronq):

The equation is 2nd order, i realized that when i applied the 2nd-order integrated rate law, and confirmed that it adhered to it. \(\dfrac{1}{[A]_t}=\dfrac{1}{[A_0]}+kt\) (2nd-order integrated rate law) \(\dfrac{1}{0.267}=\dfrac{1}{0.467}+1.6t\) \(t=1~min\) Use the same formula to find \([A]\) when \(t=13~min\).